### v1

# Hallucination Detox: Sensitive Neuron Dropout (SeND) for Large Language Model Training

Abstract
As large language models (LLMs) become increasingly deployed across various industries, concerns regarding their reliability, particularly due to hallucinations—outputs that are factually inaccurate or irrelevant to user input—have grown. Our research investigates the relationship between the training process and the emergence of hallucinations to address a key gap in existing research that focuses primarily on post hoc detection and mitigation strategies. Using models from the Pythia suite (70M–12B parameters) and several hallucination detection metrics, we analyze hallucination trends throughout training and explore LLM internal dynamics. We introduce SEnsitive Neuron Dropout (SeND), a novel training protocol designed to mitigate hallucinations by reducing variance during training. SeND achieves this by deterministically dropping neurons with significant variability on a dataset, referred to as Sensitive Neurons. In addition, we develop an unsupervised hallucination detection metric, Efficient EigenScore (EES), which approximates the traditional EigenScore in 2x speed. This efficient metric is integrated into our protocol, allowing SeND to be both computationally scalable and effective at reducing hallucinations. Our empirical evaluation demonstrates that our approach improves LLM reliability at test time by up to 40% compared to normal training while also providing an efficient method to improve factual accuracy when adapting LLMs to domains such as Wikipedia and Medical datasets.
1 Introduction
1.1 Motivation
In the age of increasingly sophisticated and ever growing Large Language Models (LLMs), we see these models being utilized in every industry imaginable. However, concerns surrounding their reliability and safety have escalated due to their exploitation and the errors that arise when they are used by the general public. One of these concerning areas discovered by the scientific community is the phenomenon of hallucinations - LLMs producing content that may not align with real-world facts, the user’s input, or training data it has seen in the past (Huang et al., 2023a). In our research, we target a specific field of hallucinations called confabulations. Confabulations are a type of hallucination where the LLM generates different responses given the same or similar inputs. This can be harmful when the generations alter between correct and factually incorrect responses.
Previous research has largely focused on identifying and addressing hallucinations in large language models (LLMs), but the impact of the training process on hallucinations remains under-explored (Huang et al., 2023a; Rawte et al., 2023; Ye et al., 2023; Hong et al., 2024; Xu et al., 2024; Chen et al., 2024; Li et al., 2024; Gao et al., 2024b). This paper addresses this gap by investigating how the iterative learning process in LLMs leads to significant variance in hallucination behavior during training. This variability indicates that the model’s factual confidence fluctuates, making it challenging to pinpoint a checkpoint at which the model has confidently learned facts.
As LLMs are deployed in high-risk industries, ensuring their reliability is crucial for user safety. However, this is not always achieved, leading to serious consequences, such as an Air Canada lawsuit over an LLM-generated incorrect policy (Garcia, 2024). Addressing such issues requires a deeper understanding of how hallucinations arise during training, enabling more reliable and efficient mitigation strategies beyond post-processing methods.
To explore these hallucination trends, we analyze models ranging from 70 million to 12 billion parameters within EleutherAI’s Pythia suite (Biderman et al., 2023), assessing them across various training checkpoints and tasks. Our goal is to validate the oscillatory behavior observed in prior studies (Li et al., 2024) through evaluation metrics from HuggingFace and EleutherAI (Hong et al., 2024; Gao et al., 2024a), and to explore the correlation between model size, training progression, and hallucination patterns.
In response to the identified variance, we introduce a novel training protocol called Sensitive Neuron Dropout (SeND). SeND is designed to emphasize confident learning of facts, and in turn reduce the likelihood of confabulations, rather than solely minimizing the stochastic gradient descent (SGD) loss (e.g., cross-entropy). By selectively dropping Sensitive Neurons—those that exhibit significant fluctuations in contextual embeddings throughout training—SeND acts as a regularization technique that reduces hallucination variance and enhances the model’s factual certainty. This provides a more reliable criterion for determining training termination, ensuring models not only achieve loss convergence but also display stable factual confidence. To maintain efficiency as model size and inference count increase, we propose the Efficient EigenScore (EES), an approximation metric for hallucination detection. EES replaces EigenScore (Chen et al., 2024), the primary metric used in our experiments, offering a scalable solution with high correlation to the original EigenScore.
Our contributions to the field can be summarized as follows, emphasizing that SeND enhances the training process but does not replace post-hoc solutions, which may still be required after training:111For the code and datasets used, refer to our GitHub repository at: https://anonymous.4open.science/r/SeND-Pythia/README.md.
1.
Empirical verification of the oscillatory nature of hallucinations in LLMs training across various model scales and detection metrics.
2.
Sensitive Neuron Dropout (SeND), a training-time method designed to reduce hallucination variance and increase model factual confidence during training.
3.
Efficient EigenScore (EES), an efficient hallucination detection metric used to keep SeND efficient, achieving up to 2x speedup with minimal effects on accuracy.
1.2 Related Work
The majority of research on hallucinations in language models has focused on detecting and mitigating this phenomenon rather than explaining its underlying causes. Recent techniques can be categorized into two main approaches: those that rely on output text or model probabilities at inference time (Manakul et al., 2023; Joshi et al., 2017; Li et al., 2023) and those that utilize internal representations or hidden layers of the model (Su et al., 2024; Chen et al., 2024; Kossen et al., 2024). While the former has demonstrated effectiveness, the latter offers deeper insights but often comes with computational trade-offs. Additionally, methods like Reinforcement Learning with Human Feedback (RLHF) have gained traction for enhancing model reliability (Yu et al., 2024). However, many of these post-hoc solutions enhance factual accuracy by layering algorithms atop pre-trained models, which can be inefficient. Our work addresses this gap by focusing on the internal dynamics of the model that contribute to hallucinations.
We use several metrics for evaluation, including Halueval (Li et al., 2023), FactScore (Min et al., 2023), SelfCheckGPT (Manakul et al., 2023), and XSum (Narayan et al., 2018), to validate our findings across different tasks. Given that the internal dynamics of the model have proven to be reliable candidates for assessing certainty and hallucination likelihood, we leverage methodologies such as EigenScore (Chen et al., 2024) and Semantic Entropy (Kossen et al., 2024), which detect hallucination risk by analyzing the variability in multiple high-temperature outputs. In our experiments, we use the EigenScore metric alongside the HELM dataset created by Su et al. (2024) to detect hallucinations during training and in the development of SeND.
Regularization techniques have been introduced to fix the issue of variability, notably random neuron dropout, used to reduce the variance and ensure that no neuron is overpowering others (Srivastava et al., 2014; Baldi & Sadowski, 2013). Work such as that done by Santra et al. (2020); Ba & Frey (2013) aims to modify random neuron dropout to change the way neurons are dropped to a more deterministic, precise manner. This has allowed the authors to drop unimportant connections in a deep neural network to ensure that class discriminative information is propagated through the model correctly (Santra et al., 2020). Inspired by this, our aim is to target hallucinatory neurons in our models to ensure that factual information is propagated through.
A significant drawback of state-of-the-art detection techniques, particularly those relying on internal model dynamics, is their efficiency. Existing methods often necessitate multiple inferences and embedding generations, making the spectral analysis of embedding matrices computationally intensive and increasingly impractical as models and datasets grow (Chen et al., 2024; Su et al., 2024). To address these challenges, we propose the use of spectral theory for efficient approximation. This approach enables scalable hallucination detection while maintaining performance. By utilizing tools such as the Density of States (DOS) and the kernel polynomial method (KPM) for approximating EigenScore (Huang et al., 2023b; Lin et al., 2014), we aim to enhance the efficiency of our analysis in the context of confabulations, which we will demonstrate empirically with EES and SeND.
2 Oscillatory Behaviour Validation
The training epochs of a transformer model can be vital in understanding the dynamics of how the model learns, particularly when trained on an unsupervised loss with stochastic gradient. Beyond the model’s architecture and the data itself, numerous factors influence the learning process such as:
Whether the loss function penalizes the learner for factual mistakes it makes or if it primarily tries to force the model to memorize the data.
While our paper does not aim to address the broader debate on whether language models truly understand language or rely heavily on memorization, our analysis of training dynamics through multiple epochs shows that reducing stochastic gradient loss does not necessarily correspond to reducing hallucinations, verifying the results Li et al. (2024) showed for the oscillatory behaviour of LLMs in hallucination during training. This highlights the need for further investigation into the relationship between optimization and factual accuracy in LLMs.
2.1 Methods
In our study, we utilize Eleuther AI’s Pythia and LMEval tools (Biderman et al., 2023; Gao et al., 2024a), to examine the development and evolution of LLMs throughout the training process. Pythia comprises a suite of 16 LLMs, all trained on public data in the same sequential order, with sizes ranging from 70 million to 12 billion parameters, 8 of which we use for our experiments. We used 20 equally spaced training checkpoints from the start to the finish for our analysis. We chose Pythia as it is based on GPT-Neo X (Black et al., 2022), which shares a similar foundational architecture to other state-of-the-art language models. Pythia’s comprehensive package of models makes it particularly suitable for our analysis, allowing us to conduct a thorough examination of the development and evolution of LLMs throughout the training process.
These models are evaluated at each checkpoint on a variety of hallucination/fact-checking metrics. To do this, we leverage the HuggingFace Hallucination Leaderboard (Hong et al., 2024), which offers comprehensive benchmarks for our experiments. There are two main components to our evaluations: Summarization and Self-consistency. For summarization, models are evaluated under the XSum dataset Narayan et al. (2018) where the model is given a dataset of BBC news articles and must give summaries of each article. A higher Rouge1 score on XSum means the data is aligning better with the provided reference summary. SelfCheckGPT (Manakul et al., 2023) is used to see if the model is uncertain with respect to a dataset of prompts for self-consistency. When the SelfCheckGPT score is high, it means the model is more likely to contradict itself on the given input and therefore, more likely to hallucinate.
2.2 How do the established iterative training processes influence LLM hallucinations?
The analysis of hallucination oscillations, as shown in Figure 1, indicates a consistent pattern across different models: oscillations persist throughout training from the initial to the final checkpoint. This finding highlights the uncertainty of halting training solely based on the convergence of training loss to its minimum to minimize hallucination. For example, in the evaluation plot of XSUM with the 12B model (1(b)), the optimal value for the hallucination metric occurs at the 60000th checkpoint, despite the model’s unsupervised loss converging by step 140000, leading to the termination of training. This evidence challenges the notion that optimizing solely for unsupervised loss in SGD guarantees learning the most accurate representation of the data. This observation is seen more drastically in 1(a), where model size has no effect on the performance of SelfCheckGPT. Instead, we observe exaggerated oscillatory behaviour within self-consistency, meaning that model size is even less effective at tackling the issue of confabulations. One potential solution to mitigate this issue could involve incorporating a regularization term into the unsupervised loss based on a hallucination detection metric discussed in Section 4.
2.3 How does model complexity affect the emergence of hallucinations throughout training?
An analysis of hallucination detection metrics reveals a diminishing rate of improvement with increased model scaling, particularly up to the 12B parameter size (Figure 1(b)). This suggests that beyond a certain point, even though there is improvement in the hallucinations, larger models do not significantly reduce hallucinations, indicating that scaling alone is not sufficient for building robust models. Instead, more refined approaches are needed to address the underlying variability in model behavior. For the following experiments, we focus on the Pythia 1B model.
3 Internal Training Dynamics
Following our investigation of the oscillatory behaviour in training, we look into the internal states of the Pythia 1B model to see what information we are able to extract. In doing so, we define a series of terms and formulas in order to understand the internal processes during the training of LLMs. This information is later used in sections 3.3 and 4 to assist us in deriving methods for improving the variance in the hallucinatory behaviour of models during training.
3.1 Sensitive Neurons
To start our analysis of the internal states, we convert the activation matrix of the model into a sentence embedding vector 3.1 which turns an activation matrix into a sentence embedding vector for input with dimension . Given its demonstrated success in hallucination detection by Su et al. (2024), we employ this sentence embedding extraction approach.
Definition 3.1 (Sentence Embedding Vector).
The Sentence Embedding Vector is a way to convert the large activation matrix into a smaller, easier to manage vector with dimension .
Where is the activation of one input , is the number of tokens in the sequence, and is the subtraction to get the penultimate layer index and the formula is adapted from Su et al. (2024). The penultimate layer of the LLM, being the layer closest to the output probabilities, is our primary focus for hallucination analysis due to its rich information about output certainty.
Next, we define the Net Change Formula 3.2 as a way to extract information from the model indicative of oscillatory behaviour between checkpoints from the sentence embedding vector.
Definition 3.2 (Net Change Formula).
Let denote the embedding of data point at neuron of the contextual embedding after checkpoint/epoch . Then we define the net change formula as
With these definitions, we can now describe the crux of our investigation: Sensitive Neurons. These Sensitive Neurons give us key parts of the model that we will prove contribute to the hallucination of LLM models. They can be used to adapt training procedures for lowering hallucination variation during training and better overall confidence at inference time. In essence, Sensitive Neurons are embedding indices in the sentence embedding from definition 3.1 that experience drastic changes between checkpoints/epochs of the training, something we believe is related to the oscillatory behaviour in hallucination performance. When finding the most Sensitive Neurons, we typically want to select the top neurons for a specific data point’s representation. In our investigation we set .
Definition 3.3 (Sensitive Neurons).
Indices of the contextual embedding for data point which exhibit the highest net change across the last checkpoints of training, indicating overall high variability during this period. This is calculated by
where is the total variability during the last checkpoints and the most Sensitive Neurons are
where N is the embedding vector size and is the desired percentile threshold.
The above definition of Sensitive Neurons is then applied to LLM hallucinations through analyses of the EigenScores. In their paper, Chen et al. (2024) define a new metric for detecting confabulations, a subclass of hallucinations. They do this by calculating an EigenScore 3.4 based on determinant calculations from multiple outputs of an LLM with a high-temperature setting (temperature set to 0.5) to encourage the LLM to produce a variety of different outputs. They propose that if an LLM is set to hallucinate on that output, the generated texts will show higher semantic variability and produce a higher EigenScore. This method achieves SOTA performance and is unsupervised as it only relies on the representations learned by the model. In the forthcoming sections, we will analyze the correlation between the EigenScore of data points during training checkpoints and the most Sensitive Neurons associated with them.
Definition 3.4 (EigenScore).
The EigenScore of data point indicates the degree of hallucination on input by the average logarithm of the eigenvalues on the covariance matrix of the multiple output generations (typically 10 in our experiments).
where denotes the eigenvalues of the regularized covariance matrix . we advise referring to Chen et al. (2024) for a more detailed analysis of this formula.
3.2 Sensitive Neuron Impact on EigenScores
To assess the correlation between Sensitive Neurons and other neurons in the embedding matrix of 10 generated outputs at a specific checkpoint, we conduct experiments aimed to determine if the presence of Sensitive Neurons indicates higher uncertainty and a greater likelihood of hallucinations.
We evaluate the Sensitive Neuron effect on the HELM dataset (Su et al., 2024), which includes outputs and internal states from six open-source LLMs based on inference over 50,000 Wikipedia articles, with human annotators labeling passages as factual or hallucinatory. To assess the impact of Sensitive Neurons on hallucination, we adapt the EigenScore method by applying it to sentence embeddings from the penultimate layer of EleutherAI’s Pythia 1B model, focusing on checkpoints between 133,000 and 143,000 training steps, where embeddings are more stable and the model has a higher degree of language understanding compared to initial checkpoints. We perform sensitive neuron dropout, removing the top 10% of Sensitive Neurons at each checkpoint, and compare the results to a baseline where 10% of neurons are randomly dropped. Additionally, we analyze the impact on hallucination-prone inputs versus non-hallucination-prone inputs to determine if Sensitive Neurons play a critical role during hallucination, without negatively affecting correct outputs.
3.2.1 What is the Effect of Sensitive Neurons on Hallucination Metrics?
Since a reduction in the EigenScore metric can be used as a proxy to show the reduction in likelihood of hallucination and an increase in the uncertainty of the model about its response, we keep using this metric in our investigations. We are able to show through our comparison of the baseline random neuron dropout and Sensitive Neuron dropout that Sensitive Neurons significantly reduce the EigenScore metric and in turn, reduce the possibility of a confabulation (Figure 2(a)) implying their highly important role in determining model’s certainty. Not only do we observe this in hallucinatory outputs, we also observe a smaller reduction in EigenScore when applying this technique to correctly answered queries (Figure 2(b)). This result indicates that our methodology has a significant effect on the uncertainty shown by an LLM. We observe that looking at the internal states of the model is an effective way to eliminate confabulating text generation in various model sizes.
3.3 Efficient EigenScore Approximation
To address the computational complexity of EigenScore calculations, particularly as LLM hidden layer sizes increase and more generations occur at higher temperatures, we develop an approximation method. This approximation, detailed in Algorithm 1, leverages the properties of Spectral Density or Density of States (DOS) to estimate EigenScore without explicitly constructing the covariance matrix. While this approximation provides a general overview of EigenScore trends, it is important to note that the output scales differ: EigenScore ranges from , whereas the approximation, referred to as Efficient EigenScore (EES), outputs values between . Since the spectrum of the matrix is altered to make EES computable and operates on its own scale, EES can be seen as a standalone metric for hallucination detection.
The computation of the Efficient EigenScore (EES) is based on two fundamental concepts: Chebyshev Polynomials and Density of States (DOS). A detailed introduction to these concepts is provided in Appendix sections B.1 and B.2. Below, we outline a brief sketch of the derivation of EES. Since Chen et al. (2024) use the covariance matrix of the embedding matrix of 10 generated sequences by the model in their methods, we represent it with and use it in our derivation.
Lemma 1.
Let . Then, for a covariance matrix with eigenvalues , we have
where are the eigenvalues of .
Proposition 1.
Note that from Proposition 1, the integral is equal to or in our application, given the integral equals , being the number of model generations.
Our objective is to simplify the integral and approximate its value, avoiding the direct computation of the covariance matrix. This approach is intended to mitigate the computational complexity and associated costs of explicitly handling the covariance matrix. Further utilizing Chebyshev Polynomials, DOS, and KPM (as introduced in Appendix B.2), we can simplify the integral mentioned in Equation 7 to , where term in DOS is approximated using Stochastic Trace Estimation and m’th Chebyshev Polynomial coefficient. Appendices B.3 and B.4 provide the derivation of this equation. Note that the simplified integral is ultimately used to approximate the EigenScore of the matrix which is ultimately equivalent to . Performance of EES approximation is closely correlated with that of the original EigenScore and can be seen to closely track the progress of EigenScore through training of Pythia 1B on the HELM dataset in Figure 7.
3.4 How does Efficient EigenScore approximation scale compared to regular EigenScore?
The efficiency of EES is compared to that of the regular EigenScore calculation with respect to scaling matrix sizes. These tests are imperative to the application of our training protocol on LLMs in Section 4, as with increasing model size, we get ever larger matrix sizes to decompose for the EigenScore calculation. To elucidate these scaling effects, we conduct a grid search over two important parameters: Matrix size (Figure 3) and Moments used for EES calculation (Figure B.5). The difference between EES time in comparison to EigenScore when increasing the number of columns and rows is visualized in Figure 3 using a moments value of 20. It is evident that EES provides a significant computational advantage when increasing the number of columns or rows. Remarkably, at matrix size , EES nearly halves the computation time of regular EigenScore calculation at around 4 seconds whereas EigenScore takes approximately 7 seconds to calculate. We can then deduce that given a good enough approximation, EES provides a significant reduction in computational complexity as model and matrix size increase.
4 SEnsitive Neuron Dropout (SeND)
Building on the findings from Section 3.2, and aiming to reduce variance in the factual uncertainty of LLMs during training, this section introduces SeND, an efficient and transferable framework for training LLMs. SeND integrates the EES method discussed in Section 3.3 to enhance computational efficiency while addressing variance in sensitive neuron behavior. By identifying sensitive neurons, which contribute to the oscillatory behavior of hallucinations during training, SeND deterministically drops these neurons based on a small subset of the training data. This approach ensures an increase in the model’s factual certainty by the end of training as explained in Algorithm 2.
4.1 SeND Experiment Setup
To evaluate SeND, we use Eleuther AI’s Pythia 1B model, continuing its training on specific datasets rather than restarting pretraining to maintain efficiency. We continue the training of the fully trained model on two datasets: HELM, consisting of Wikipedia text (Su et al., 2024), and MedHALT, a medical dataset emulating real-world entrance exam questions (Pal et al., 2023). Due to the importance of factual accuracy in the medical domain, MedHALT was chosen to assess SeND’s impact on hallucination mitigation. Both datasets were tested in two sizes: 200 and 2,000 points (referred to as 2k). SeND implements the EigenScore reduction technique from Section 3.2 and detects Sensitive Neurons using a 3-epoch window on a specialized hallucination tracking dataset. Sensitive Neurons in the penultimate layer are identified based on their variability across epochs and are deterministically dropped for the subsequent 3 training epochs. This dropout process is repeated at each 3-epoch interval until the training loss converges, effectively mitigating hallucination tendencies and refining the model. While we acknowledge that the number of subsequent epochs for dropout could be adjusted, we limit our approach to 3 for ease of computation.
4.2 Performance of SeND on Pythia 1B
The results of training Pythia 1B on HELM and MedHALT 200 are illustrated in Figure 4. To validate that the EES method accurately approximates the EigenScore metric; we compare the model’s progress during training (up to loss convergence) and assess whether the resulting graphs are similar. These results are detailed in Appendix B.6. Upon confirming that EES provides a reliable approximation of the EigenScore, we proceed to compare the performance of Pythia 1B trained using standard training with SeND on HELM and MedHALT 2k (Figure 4). In the case of training on HELM with the regular protocol, we observe results consistent with previous findings: while the model successfully reduces loss, it fails to optimize for hallucination, as evidenced by the increasing EES metric (green line in Figure 4(a)). Conversely, training with SeND reveals a consistent trajectory toward reducing both EES and loss, as depicted by the blue line.
To assess the effectiveness of SeND in comparison to other state-of-the-art factuality metrics, we employ the FactScore metric from Min et al. (2023), which quantifies the factual accuracy of content generated by large language models (LLMs). The fact-checking is conducted using the HELM dataset on the final trained models (1B SeND and 1B Normal Training). A higher FactScore indicates improved factual precision. When evaluated on 100 data points from the HELM dataset, the 1B SeND model achieves a FactScore of 0.07, whereas the 1B Normal Training model attains 0.05, demonstrating a 40% improvement in factual accuracy, even during test time. This highlights the efficacy of SeND in enhancing the factual certainty of the model. Note that SeND is not a replacement for post-hoc methods like RAG (Gao et al., 2024b), but rather to complement them.
The robustness of this training protocol is essential as we aim for it to be applied across many different fields. In light of this, we present the results of finetuning on MedHALT 2k with and without SeND in Figure 4(b). We observe a similar trend in 4(b) as shown in Figure 4(a), where standard finetuning increases the EES score throughout training, showing that the model is not taking into account hallucinations and factuality during its training. In Figure 4(b), there is an improvement in the trajectory of EES as training continues, showing that our model is in fact able to incorporate factuality as a metric to account for during training of the model. Although the decrease in EES is marginal for MedHALT 2k, we can attribute this to the dataset it is being trained on. The relatively small difference observed between the training protocol behaviours could be due to MedHALT 2k data never being seen before the finetuning phase whereas HELM data has been seen. In this case, it may be beneficial to delay the onset of SeND, as high variability between checkpoints on new training instances is expected.
5 Conclusion & Future Work
In this paper, we presented a protocol to refine the current training methods of LLMs based on experiments showing oscillatory behaviour with respect to hallucinations throughout training (Figure 1). To do this we used the internal states of LLMs, specifically the penultimate layer activations during inference on a specialized dataset. We present an initial method of reducing hallucinations based on the principles of EigenScore metrics introduced by Chen et al. (2024). We showed empirically that our Sensitive Neuron detection method significantly reduces the EigenScore on inference of LLMs throughout various stages of training (Figure 2). Following the success of the Sensitive Neuron method, we moved on to the application of a hallucination reduction method on training. We show through finetuning that we are able to fix the oscillatory behaviour initially seen throughout training and reduce the EES of finetuned models as shown in Figure 4 by modifying the internal mechanics of training with SEnsitive Neuron Dropout. At test time we achieve a 40% increase in FactScore performance, verifying that SeND provides a substantial improvement to current training protocols.
To further advance our work, we plan to scale SeND to larger datasets and models, as current experiments were limited by compute constraints with larger LLMs. Demonstrating SeND’s effectiveness on larger open-source models like Meta’s LLaMA 3.1 (Dubey et al., 2024) will provide crucial evidence for organizations developing state-of-the-art LLMs to incorporate SeND into their training protocols, ultimately improving model safety. Given that SeND targets variance reduction during training, we anticipate even greater gains on larger LLMs, where the higher inherent variance may amplify the regularization effect and lead to more significant improvements.
Appendix A Additional Experiments
A.1 Drastic Embedding Changes leading to Sensitive Neurons
Looking at internal states of the model allows getting a deeper understanding of the dynamics that could be leading to the oscillatory behaviour seen in Figure 1. To do this, we record the net change (Definition 3.2) between checkpoints of the penultimate layer where one checkpoint would be the correct answer and the next would hallucinate. This net change with respect to various different input texts is plotted in Figure 5. It can be observed that there were specific embedding activations that experienced drastically more change relative to the rest of the embeddings. This is the main source of motivation to further define Sensitive Neurons (Definition 3.3).
Appendix B Efficient EigenScore (EES) Derivation
B.1 Background: Chebyshev polynomials
Chebyshev polynomials are a sequence of orthogonal polynomials in the interval – orthogonality property shown in equation 8 – that are widely used in numerical analysis, approximation theory, and other areas of applied mathematics. In this work, we are mainly concerned with the Chebyshev polynomials of the first kind with the recurrence relation shown in equation 9. Note that this recurrence could also be applied to matrices. Any function defined in the interval can be approximated with the Chebyshev expansion as shown in 10.
B.2 Background: DOS and KPM
Let be a symmetric matrix with an eigendecomposition , where and is orthogonal. The spectral density induced by is the generalized function:
where is the Dirac delta function. For any analytic test function , the integral of with respect to is:
Dong et al. (2019) introduced KPM as a numerical technique to approximate DOS. KPM approximates DOS by expanding it in terms of chebyshev polynomials. Requiring the matrix’s spectrum to be supported in the interval , KPM approximates DOS with the following formula, being the eigen value of the matrix and approximated by Stochastic Trace Estimation:
In the application for hallucination detection, we can use equation 14 to derive a formula for the EigenScore approximation using the properties of Chebyshev polynomials and DOS.
B.3 Stochastic Trace Estimation on Embedding Matrix
We are interested in computing the term of DOS relying solely on the embedding matrix therefore we need to rewrite as follows:
where can be computed using the Chebyshev polynomials of matrix .
Each term can be computed with a matrix-vector multiplication.
B.4 EES Integral Calculation
Given the orthogonality of the Chebyshev polynomials, we can simplify the integral mentioned in proposition 1. To approximate the EigenScore, we will expand in terms of Chebyshev polynomials and use their orthogonality to simplify the integral.
Expanding and Integrating
To approximate the integral:
Substitute the Chebyshev Expansion for DOS:
where:
Distribute in the integral:
Evaluate the Integral Using Orthogonality:
To simplify the integral,
First, express as a series of Chebyshev polynomials:
Then:
Note: The lower bound of the integral is 0 as the matrix is defined in the spectrum .
Using the orthogonality, we get:
So the integral simplifies to:
B.5 Efficient EigenScore Moments
Figure B.5 presents the effect of using different moment values as the number of matrix rows increases with respect to time. This is an important hyperparameter to tune as increasing the number of moments on EES correlates to having a more accurate and representative approximation of the EigenScore. We observe that as moments in EES increase, the time to calculate EES increases. From this result, we conclude that selecting a moment value of under 50 would provide a balanced trade-off between accuracy and calculation time.
Effect of changing number of moments on EES calculation time (seconds). More moments gives more accurate approximation but higher computation time.
B.6 EigenScore and EES training trajectories
To demonstrate that our EigenScore approximation method, EES, is a good metric, we record the progress of Pythia 1B finetuning on the HELM dataset using both EigenScore and EES hallucination performance metrics (Figure 7). Albeit a different scale and window, the trajectories, magnitude and shape of the graphs are nearly identical while EES takes only 4 minutes to calculate and EigenScore takes approximately 8, an astounding 2x increase in compute speed. These results show that our metric closely resembles the target metric while greatly reducing the required computational resources.

### v2

@@
-Hallucination Detox: Sensitive Neuron Dropout (SeND) for Large Language Model Training
+Hallucination Detox: Sensitivity Dropout ( SenD) for Large Language Model Training
 Abstract
-As large language models (LLMs) become increasingly deployed across various industries, concerns regarding their reliability, particularly due to hallucinations—outputs that are factually inaccurate or irrelevant to user input—have grown.
+As large language models (LLMs) are increasingly deployed across various industries, concerns regarding their reliability, particularly due to hallucinations—outputs that are factually inaccurate or irrelevant to user input—have grown.
 Our research investigates the relationship between the training process and the emergence of hallucinations to address a key gap in existing research that focuses primarily on post hoc detection and mitigation strategies.
 Using models from the Pythia suite (70M–12B parameters) and several hallucination detection metrics, we analyze hallucination trends throughout training and explore LLM internal dynamics.
-We introduce SEnsitive Neuron Dropout (SeND), a novel training protocol designed to mitigate hallucinations by reducing variance during training.
-SeND achieves this by deterministically dropping neurons with significant variability on a dataset, referred to as Sensitive Neurons.
+We introduce Sensitivity Dropout ( SenD), a novel training protocol designed to mitigate hallucinations by reducing variance during training.
+SenD achieves this by deterministically dropping embedding indices with significant variability, referred to as Sensitive Embedding Indices.
 In addition, we develop an unsupervised hallucination detection metric, Efficient EigenScore (EES), which approximates the traditional EigenScore in 2x speed.
 This efficient metric is integrated into our protocol, allowing SenD to be both computationally scalable and effective at reducing hallucinations.
-Our empirical evaluation demonstrates that our approach improves LLM reliability at test time by up to 40% compared to normal training while also providing an efficient method to improve factual accuracy when adapting LLMs to domains such as Wikipedia and Medical datasets.
+Our empirical evaluation demonstrates that our approach improves LLM reliability at test time by up to 40% compared to normal training while also providing an efficient method to improve factual accuracy when adapting LLMs to Wikipedia, Medical, and LegalBench domains.
 1 Introduction
 1.1 Motivation
-In the age of increasingly sophisticated and ever growing Large Language Models (LLMs), we see these models being utilized in every industry imaginable.
-However, concerns surrounding their reliability and safety have escalated due to their exploitation and the errors that arise when they are used by the general public.
+As Large Language Models (LLMs) become more sophisticated and widespread across industries, concerns about their reliability and safety have grown due to misuse and user errors.
 One of these concerning areas discovered by the scientific community is the phenomenon of hallucinations - LLMs producing content that may not align with real-world facts, the user’s input, or training data it has seen in the past (Huang et al., 2023a).
-In our research, we target a specific field of hallucinations called confabulations.
-Confabulations are a type of hallucination where the LLM generates different responses given the same or similar inputs.
+In our research, we target a specific field of hallucinations called confabulations which occur when the LLM generates different responses given the same or similar inputs.
 This can be harmful when the generations alter between correct and factually incorrect responses.
@@
 Addressing such issues requires a deeper understanding of how hallucinations arise during training, enabling more reliable and efficient mitigation strategies beyond post-processing methods.
-To explore these hallucination trends, we analyze models ranging from 70 million to 12 billion parameters within EleutherAI’s Pythia suite (Biderman et al., 2023), assessing them across various training checkpoints and tasks.
-Our goal is to validate the oscillatory behavior observed in prior studies (Li et al., 2024) through evaluation metrics from HuggingFace and EleutherAI (Hong et al., 2024; Gao et al., 2024a), and to explore the correlation between model size, training progression, and hallucination patterns.
-In response to the identified variance, we introduce a novel training protocol called Sensitive Neuron Dropout (SeND).
-SeND is designed to emphasize confident learning of facts, and in turn reduce the likelihood of confabulations, rather than solely minimizing the stochastic gradient descent (SGD) loss (e.g., cross-entropy).
-By selectively dropping Sensitive Neurons—those that exhibit significant fluctuations in contextual embeddings throughout training—SeND acts as a regularization technique that reduces hallucination variance and enhances the model’s factual certainty.
+To explore these hallucination trends, we analyze models ranging from 70 million to 12 billion parameters within Pythia suite (Biderman et al., 2023), assessing them across various training checkpoints and tasks.
+Our goal is to validate the oscillatory behavior observed by Li et al.
+(2024) through evaluation metrics including HaluEval (Li et al., 2023), FactScore (Min et al., 2023), SelfCheckGPT (Manakul et al., 2023), and XSum (Narayan et al., 2018).
+Utilizing the reliability of internal model dynamics for quantifying hallucination likelihood, we use EigenScore (Chen et al., 2024) and Semantic Entropy (Kossen et al., 2024) to detect hallucination risk by analyzing variability in high-temperature outputs.
+Experiments utilize EigenScore and the HELM dataset (Su et al., 2024) to identify hallucinations during training.
+In response to this variance, we introduce a novel training protocol called Sensitivity Dropout ( SenD).
+SenD is designed to emphasize confident learning of facts, and in turn reduce the likelihood of confabulations, rather than solely minimizing the training loss.
+By selectively dropping Sensitive Embedding Indices—those exhibiting significant fluctuations throughout training— SenD acts as a technique that reduces hallucination variance and enhances the model’s factual certainty.
 This provides a more reliable criterion for determining training termination, ensuring models not only achieve loss convergence but also display stable factual confidence.
-To maintain efficiency as model size and inference count increase, we propose the Efficient EigenScore (EES), an approximation metric for hallucination detection.
-EES replaces EigenScore (Chen et al., 2024), the primary metric used in our experiments, offering a scalable solution with high correlation to the original EigenScore.
+To maintain efficiency as model size and inference count increase, we propose the Efficient EigenScore (EES), a novel metric for hallucination detection.
+EES replaces EigenScore (Chen et al., 2024), the primary metric used in our experiments, offering a scalable solution with high correlation to EigenScore.
 Our contributions to the field can be summarized as follows, emphasizing that SenD enhances the training process but does not replace post-hoc solutions, which may still be required after training:111For the code and datasets used, refer to our GitHub repository at: https://anonymous.4open.science/r/SeND-Pythia/README.md.
@@
 2.
-Sensitive Neuron Dropout (SeND), a training-time method designed to reduce hallucination variance and increase model factual confidence during training.
+Sensitivity Dropout ( SenD), a training-time method designed to reduce hallucination variance and increase model factual confidence during training.
 3.
@@
 The majority of research on hallucinations in language models has focused on detecting and mitigating this phenomenon rather than explaining its underlying causes.
-Recent techniques can be categorized into two main approaches: those that rely on output text or model probabilities at inference time (Manakul et al., 2023; Joshi et al., 2017; Li et al., 2023) and those that utilize internal representations or hidden layers of the model (Su et al., 2024; Chen et al., 2024; Kossen et al., 2024).
+Recent techniques can be categorized into two main approaches: those that rely on output probabilities at inference time (Manakul et al., 2023; Joshi et al., 2017; Li et al., 2023) and those that utilize internal representations or hidden layers of the model (Su et al., 2024; Chen et al., 2024; Kossen et al., 2024).
 While the former has demonstrated effectiveness, the latter offers deeper insights but often comes with computational trade-offs.
@@
 Our work addresses this gap by focusing on the internal dynamics of the model that contribute to hallucinations.
-We use several metrics for evaluation, including Halueval (Li et al., 2023), FactScore (Min et al., 2023), SelfCheckGPT (Manakul et al., 2023), and XSum (Narayan et al., 2018), to validate our findings across different tasks.
-Given that the internal dynamics of the model have proven to be reliable candidates for assessing certainty and hallucination likelihood, we leverage methodologies such as EigenScore (Chen et al., 2024) and Semantic Entropy (Kossen et al., 2024), which detect hallucination risk by analyzing the variability in multiple high-temperature outputs.
-In our experiments, we use the EigenScore metric alongside the HELM dataset created by Su et al.
-(2024) to detect hallucinations during training and in the development of SeND.
 Regularization techniques have been introduced to fix the issue of variability, notably random neuron dropout, used to reduce the variance and ensure that no neuron is overpowering others (Srivastava et al., 2014; Baldi & Sadowski, 2013).
@@
 This has allowed the authors to drop unimportant connections in a deep neural network to ensure that class discriminative information is propagated through the model correctly (Santra et al., 2020).
-Inspired by this, our aim is to target hallucinatory neurons in our models to ensure that factual information is propagated through.
-A significant drawback of state-of-the-art detection techniques, particularly those relying on internal model dynamics, is their efficiency.
-Existing methods often necessitate multiple inferences and embedding generations, making the spectral analysis of embedding matrices computationally intensive and increasingly impractical as models and datasets grow (Chen et al., 2024; Su et al., 2024).
-To address these challenges, we propose the use of spectral theory for efficient approximation.
-This approach enables scalable hallucination detection while maintaining performance.
-By utilizing tools such as the Density of States (DOS) and the kernel polynomial method (KPM) for approximating EigenScore (Huang et al., 2023b; Lin et al., 2014), we aim to enhance the efficiency of our analysis in the context of confabulations, which we will demonstrate empirically with EES and SeND.
+Inspired by this, our aim is to target hallucinatory embedding indices in our models to ensure that factual information is propagated through.
+State-of-the-art hallucination metrics, especially those based on internal model dynamics, rely on spectral analysis and embedding matrix computations.
+Methods like EigenScore (Chen et al., 2024) and Semantic Entropy (Kossen et al., 2024) effectively assess hallucination risk but require multiple inferences, making them computationally demanding as models scale.
+Tools such as the Density of States (DOS) and the kernel polynomial method (KPM) have been explored to approximate spectral properties efficiently (Huang et al., 2023b; Lin et al., 2014).
+Building on these advancements, our work integrates efficient spectral analysis methods into hallucination detection, demonstrated through EES and SenD.
 2 Oscillatory Behaviour Validation
-The training epochs of a transformer model can be vital in understanding the dynamics of how the model learns, particularly when trained on an unsupervised loss with stochastic gradient.
-Beyond the model’s architecture and the data itself, numerous factors influence the learning process such as:
-Whether the loss function penalizes the learner for factual mistakes it makes or if it primarily tries to force the model to memorize the data.
-While our paper does not aim to address the broader debate on whether language models truly understand language or rely heavily on memorization, our analysis of training dynamics through multiple epochs shows that reducing stochastic gradient loss does not necessarily correspond to reducing hallucinations, verifying the results Li et al.
-(2024) showed for the oscillatory behaviour of LLMs in hallucination during training.
-This highlights the need for further investigation into the relationship between optimization and factual accuracy in LLMs.
-2.1 Methods
-In our study, we utilize Eleuther AI’s Pythia and LMEval tools (Biderman et al., 2023; Gao et al., 2024a), to examine the development and evolution of LLMs throughout the training process.
-Pythia comprises a suite of 16 LLMs, all trained on public data in the same sequential order, with sizes ranging from 70 million to 12 billion parameters, 8 of which we use for our experiments.
-We used 20 equally spaced training checkpoints from the start to the finish for our analysis.
-We chose Pythia as it is based on GPT-Neo X (Black et al., 2022), which shares a similar foundational architecture to other state-of-the-art language models.
-Pythia’s comprehensive package of models makes it particularly suitable for our analysis, allowing us to conduct a thorough examination of the development and evolution of LLMs throughout the training process.
-These models are evaluated at each checkpoint on a variety of hallucination/fact-checking metrics.
-To do this, we leverage the HuggingFace Hallucination Leaderboard (Hong et al., 2024), which offers comprehensive benchmarks for our experiments.
-There are two main components to our evaluations: Summarization and Self-consistency.
-For summarization, models are evaluated under the XSum dataset Narayan et al.
-(2018) where the model is given a dataset of BBC news articles and must give summaries of each article.
-A higher Rouge1 score on XSum means the data is aligning better with the provided reference summary.
-SelfCheckGPT (Manakul et al., 2023) is used to see if the model is uncertain with respect to a dataset of prompts for self-consistency.
-When the SelfCheckGPT score is high, it means the model is more likely to contradict itself on the given input and therefore, more likely to hallucinate.
+The training checkpoints of a transformer model can be vital in understanding the dynamics of how the model learns.
+Beyond the model’s architecture and the data itself, numerous factors influence the learning process such as: whether the loss function penalizes the learner for factual mistakes it makes or if it primarily tries to force the model to memorize the data.
+While our paper does not aim to address the broader debate on whether LLMs truly understand language or rely on memorization, our analysis of training dynamics through multiple checkpoints shows that converging the training loss does not necessarily correspond to reducing hallucinations, verifying the results by Li et al.
+(2024) for LLM oscillatory hallucination behaviour during training.
+Our study utilizes Eleuther AI’s Pythia and LMEval tools (Biderman et al., 2023; Gao et al., 2024a) to examine the development and evolution of LLMs throughout the training process.
+Pythia comprises a suite of 16 LLMs, all trained on public data in the same sequential order, with sizes ranging from 70 million to 12 billion parameters.
+We use 20 equally spaced training checkpoints from the start to the finish for our analysis.
+These models are evaluated at each checkpoint on a variety of hallucination metrics such as SelfCheckGPT for Self-consistency (Manakul et al., 2023), XSum for Summarization (Narayan et al., 2018), Perplexity, and HaluEval (Li et al., 2023) for Question Answering (QA) tasks.
+High SelfCheckGPT Score means that the model is more likely to contradict itself on the given input, a higher Rouge1 score on XSum means the data is aligning better with the provided reference summary, lower perplexity implies higher prediction confidence, and higher Exact Match, Accuracy, and Correctness implies better performance of the model on QA tasks.
 2.1 How do the established iterative training processes influence LLM hallucinations?
 The analysis of hallucination oscillations, as shown in Figure 1, indicates a consistent pattern across different models: oscillations persist throughout training from the initial to the final checkpoint.
-This finding highlights the uncertainty of halting training solely based on the convergence of training loss to its minimum to minimize hallucination.
-For example, in the evaluation plot of XSUM with the 12B model (1(b)), the optimal value for the hallucination metric occurs at the 60000th checkpoint, despite the model’s unsupervised loss converging by step 140000, leading to the termination of training.
+This finding highlights the uncertainty of halting training solely based on the convergence of training loss.
+For instance, in QA settings, the optimal Exact Match of the outputs with ground truths is achieved in earlier checkpoints.
 This evidence challenges the notion that optimizing solely for unsupervised loss in SGD guarantees learning the most accurate representation of the data.
-This observation is seen more drastically in 1(a), where model size has no effect on the performance of SelfCheckGPT.
-Instead, we observe exaggerated oscillatory behaviour within self-consistency, meaning that model size is even less effective at tackling the issue of confabulations.
-One potential solution to mitigate this issue could involve incorporating a regularization term into the unsupervised loss based on a hallucination detection metric discussed in Section 4.
+This observation is seen more drastically in 1(a), where model size has nearly no effect on the performance of SelfCheckGPT.
+Instead, we observe oscillatory behaviour within self-consistency, implying that model size is not much effective at tackling the issue of confabulations verified by results in Appendix A.2 as well.
+Our mitigation approach ( SenD) is discussed in Section 4.
 2.2 How does model complexity affect the emergence of hallucinations throughout training?
-An analysis of hallucination detection metrics reveals a diminishing rate of improvement with increased model scaling, particularly up to the 12B parameter size (Figure 1(b)).
+An analysis of hallucination detection metrics reveals a diminishing rate of improvement with increased model scaling, particularly up to the 12B parameter size (Appendix A.2 for the study).
 This suggests that beyond a certain point, even though there is improvement in the hallucinations, larger models do not significantly reduce hallucinations, indicating that scaling alone is not sufficient for building robust models.
@@
 This information is later used in sections 3.3 and 4 to assist us in deriving methods for improving the variance in the hallucinatory behaviour of models during training.
-3.1 Sensitive Neurons
+3.1 Sensitive Embedding Indices
 To start our analysis of the internal states, we convert the activation matrix of the model into a sentence embedding vector 3.1 which turns an activation matrix into a sentence embedding vector for input with dimension .
@@
 The Sentence Embedding Vector is a way to convert the large activation matrix into a smaller, easier to manage vector with dimension .
-Where is the activation of one input , is the number of tokens in the sequence, and is the subtraction to get the penultimate layer index and the formula is adapted from Su et al.
+Where is the activation of one input , is the number of tokens in the sequence, is the token embedding activation matrix, and is the subtraction to get the penultimate layer index and the formula is adapted from Su et al.
 (2024).
@@
 Definition 3.2 (Net Change Formula).
-Let denote the embedding of data point at neuron of the contextual embedding after checkpoint/epoch . Then we define the net change formula as
-With these definitions, we can now describe the crux of our investigation: Sensitive Neurons.
-These Sensitive Neurons give us key parts of the model that we will prove contribute to the hallucination of LLM models.
+Let denote the embedding of data point at embedding index of the contextual embedding after checkpoint . Then we define the net change formula as
+With these definitions, we can now describe the crux of our investigation: Sensitive Embedding Indices (SEIs).
+These SEIs give us key parts of the model that we will prove contribute to the hallucination of LLM models.
 They can be used to adapt training procedures for lowering hallucination variation during training and better overall confidence at inference time.
-In essence, Sensitive Neurons are embedding indices in the sentence embedding from definition 3.1 that experience drastic changes between checkpoints/epochs of the training, something we believe is related to the oscillatory behaviour in hallucination performance.
-When finding the most Sensitive Neurons, we typically want to select the top neurons for a specific data point’s representation.
+In essence, SEIs are embedding indices in the sentence embedding from definition 3.1 that experience drastic changes between checkpoints of the training, something we believe is related to the oscillatory behaviour in hallucination performance.
+When finding the most sensitive embedding indices, we typically want to select the top embedding indices for a specific data point’s representation.
 In our investigation we set .
-Definition 3.3 (Sensitive Neurons).
+Definition 3.3 ( Sensitive Embedding Indices - SEIs).
 Indices of the contextual embedding for data point which exhibit the highest net change across the last checkpoints of training, indicating overall high variability during this period.
 This is calculated by
-where is the total variability during the last checkpoints and the most Sensitive Neurons are
+where is the total variability during the last checkpoints and the most sensitive embedding indices are
 where N is the embedding vector size and is the desired percentile threshold.
-The above definition of Sensitive Neurons is then applied to LLM hallucinations through analyses of the EigenScores.
+The above definition of SEIs is then applied to LLM hallucinations through analyses of the EigenScores.
 In their paper, Chen et al.
@@
 This method achieves SOTA performance and is unsupervised as it only relies on the representations learned by the model.
-In the forthcoming sections, we will analyze the correlation between the EigenScore of data points during training checkpoints and the most Sensitive Neurons associated with them.
+In the forthcoming sections, we will analyze the correlation between the EigenScore of data points during training checkpoints and the most sensitive embedding indices associated with them.
 Definition 3.4 (EigenScore).
@@
 where denotes the eigenvalues of the regularized covariance matrix . we advise referring to Chen et al. (2024) for a more detailed analysis of this formula.
-3.2 Sensitive Neuron Impact on EigenScores
-To assess the correlation between Sensitive Neurons and other neurons in the embedding matrix of 10 generated outputs at a specific checkpoint, we conduct experiments aimed to determine if the presence of Sensitive Neurons indicates higher uncertainty and a greater likelihood of hallucinations.
-We evaluate the Sensitive Neuron effect on the HELM dataset (Su et al., 2024), which includes outputs and internal states from six open-source LLMs based on inference over 50,000 Wikipedia articles, with human annotators labeling passages as factual or hallucinatory.
-To assess the impact of Sensitive Neurons on hallucination, we adapt the EigenScore method by applying it to sentence embeddings from the penultimate layer of EleutherAI’s Pythia 1B model, focusing on checkpoints between 133,000 and 143,000 training steps, where embeddings are more stable and the model has a higher degree of language understanding compared to initial checkpoints.
-We perform sensitive neuron dropout, removing the top 10% of Sensitive Neurons at each checkpoint, and compare the results to a baseline where 10% of neurons are randomly dropped.
-Additionally, we analyze the impact on hallucination-prone inputs versus non-hallucination-prone inputs to determine if Sensitive Neurons play a critical role during hallucination, without negatively affecting correct outputs.
-3.2.1 What is the Effect of Sensitive Neurons on Hallucination Metrics?
-Since a reduction in the EigenScore metric can be used as a proxy to show the reduction in likelihood of hallucination and an increase in the uncertainty of the model about its response, we keep using this metric in our investigations.
-We are able to show through our comparison of the baseline random neuron dropout and Sensitive Neuron dropout that Sensitive Neurons significantly reduce the EigenScore metric and in turn, reduce the possibility of a confabulation (Figure 2(a)) implying their highly important role in determining model’s certainty.
+3.2 Sensitive Embedding Index Impact on EigenScores
+To assess the correlation between SEIs and other indices in the embedding matrix of 10 generated outputs at a specific checkpoint, we conduct experiments aimed to determine if the presence of SEIs indicates higher uncertainty and a greater likelihood of hallucinations.
+We evaluate the SEI effect on the HELM dataset (Su et al., 2024), which includes outputs from six open-source LLMs based on inference over 50,000 Wikipedia articles, with human annotators labeling passages as factual or hallucinatory.
+This dataset was selected as Wikipedia is one of the main fact sources people refer to, therefore LLMs should be robust to this type of information as well as it being created with the intended use case of hallucination detection.
+To assess the impact of SEIs on hallucination, we adapt the EigenScore method by applying it to sentence embeddings from the penultimate layer of EleutherAI’s Pythia 1B model, focusing on checkpoints between 133,000 and 143,000 training steps, where embeddings are more stable and the model has a higher degree of language understanding compared to initial checkpoints.
+We perform SEI dropout, dropping the top 10% of SEIs at each checkpoint, and compare the results to a baseline where 10% of embedding indices are randomly dropped.
+Additionally, we analyze the impact on hallucination-prone inputs versus non-hallucination-prone inputs to determine if SEIs play a critical role during hallucination, without negatively affecting correct outputs.
+3.2.1 What is the Effect of Sensitive Embedding Indices on Hallucination Metrics?
+Since a reduction in the EigenScore metric can be used as a proxy to show the reduction in likelihood of hallucination, we keep using this metric in our investigations.
+We are able to show through our comparison of the baseline random embedding index dropout and SEI dropout that SEIs significantly reduce the EigenScore metric and in turn, reduce the possibility of a confabulation (Figure 2(a)), implying their highly important role in determining model’s certainty.
 Not only do we observe this in hallucinatory outputs, we also observe a smaller reduction in EigenScore when applying this technique to correctly answered queries (Figure 2(b)).
 This result indicates that our methodology has a significant effect on the uncertainty shown by an LLM.
-We observe that looking at the internal states of the model is an effective way to eliminate confabulating text generation in various model sizes.
+We observe that the internal states of the model are effective in the elimination of confabulating text generation in various model sizes.
 3.3 Efficient EigenScore Approximation
-To address the computational complexity of EigenScore calculations, particularly as LLM hidden layer sizes increase and more generations occur at higher temperatures, we develop an approximation method.
+To address the computational complexity of EigenScore calculations, particularly as LLM hidden layer sizes increase, we develop an approximation method.
 This approximation, detailed in Algorithm 1, leverages the properties of Spectral Density or Density of States (DOS) to estimate EigenScore without explicitly constructing the covariance matrix.
@@
 Note that the simplified integral is ultimately used to approximate the EigenScore of the matrix which is ultimately equivalent to .
-Performance of EES approximation is closely correlated with that of the original EigenScore and can be seen to closely track the progress of EigenScore through training of Pythia 1B on the HELM dataset in Figure 7.
-3.4 How does Efficient EigenScore approximation scale compared to regular EigenScore?
+Performance of EES approximation is closely correlated with that of the original EigenScore which can be seen in Figure 10.
+3.4 How does EES scale compared to regular EigenScore?
 The efficiency of EES is compared to that of the regular EigenScore calculation with respect to scaling matrix sizes.
-These tests are imperative to the application of our training protocol on LLMs in Section 4, as with increasing model size, we get ever larger matrix sizes to decompose for the EigenScore calculation.
-To elucidate these scaling effects, we conduct a grid search over two important parameters: Matrix size (Figure 3) and Moments used for EES calculation (Figure B.5).
+These tests are imperative to the application of our training protocol on increasing LLM sizes in Section 4 due to larger matrix sizes to decompose for the EigenScore calculation.
+We conduct a grid search over two important parameters: Matrix size (Figure 3) and Moments used for EES calculation (Figure B.5).
 The difference between EES time in comparison to EigenScore when increasing the number of columns and rows is visualized in Figure 3 using a moments value of 20.
@@
 We can then deduce that given a good enough approximation, EES provides a significant reduction in computational complexity as model and matrix size increase.
-4 SEnsitive Neuron Dropout (SeND)
+4 Sensitivity Dropout (SenD)
 Building on the findings from Section 3.2, and aiming to reduce variance in the factual uncertainty of LLMs during training, this section introduces SenD, an efficient and transferable framework for training LLMs.
-SeND integrates the EES method discussed in Section 3.3 to enhance computational efficiency while addressing variance in sensitive neuron behavior.
-By identifying sensitive neurons, which contribute to the oscillatory behavior of hallucinations during training, SeND deterministically drops these neurons based on a small subset of the training data.
+SenD integrates the EES method discussed in Section 3.3 to enhance computational efficiency while addressing variance in SEI behavior.
+By identifying SEIs, which contribute to the oscillatory behavior of hallucinations during training, SenD deterministically drops these embedding indices based on a small subset of the training data.
 This approach ensures an increase in the model’s factual certainty by the end of training as explained in Algorithm 2.
 4.1 SenD Experiment Setup
-To evaluate SeND, we use Eleuther AI’s Pythia 1B model, continuing its training on specific datasets rather than restarting pretraining to maintain efficiency.
-We continue the training of the fully trained model on two datasets: HELM, consisting of Wikipedia text (Su et al., 2024), and MedHALT, a medical dataset emulating real-world entrance exam questions (Pal et al., 2023).
-Due to the importance of factual accuracy in the medical domain, MedHALT was chosen to assess SeND’s impact on hallucination mitigation.
-Both datasets were tested in two sizes: 200 and 2,000 points (referred to as 2k).
-SeND implements the EigenScore reduction technique from Section 3.2 and detects Sensitive Neurons using a 3-epoch window on a specialized hallucination tracking dataset.
-Sensitive Neurons in the penultimate layer are identified based on their variability across epochs and are deterministically dropped for the subsequent 3 training epochs.
-This dropout process is repeated at each 3-epoch interval until the training loss converges, effectively mitigating hallucination tendencies and refining the model.
-While we acknowledge that the number of subsequent epochs for dropout could be adjusted, we limit our approach to 3 for ease of computation.
-4.2 Performance of SeND on Pythia 1B
-The results of training Pythia 1B on HELM and MedHALT 200 are illustrated in Figure 4.
-To validate that the EES method accurately approximates the EigenScore metric; we compare the model’s progress during training (up to loss convergence) and assess whether the resulting graphs are similar.
-These results are detailed in Appendix B.6.
-Upon confirming that EES provides a reliable approximation of the EigenScore, we proceed to compare the performance of Pythia 1B trained using standard training with SeND on HELM and MedHALT 2k (Figure 4).
-In the case of training on HELM with the regular protocol, we observe results consistent with previous findings: while the model successfully reduces loss, it fails to optimize for hallucination, as evidenced by the increasing EES metric (green line in Figure 4(a)).
-Conversely, training with SeND reveals a consistent trajectory toward reducing both EES and loss, as depicted by the blue line.
-To assess the effectiveness of SeND in comparison to other state-of-the-art factuality metrics, we employ the FactScore metric from Min et al.
-(2023), which quantifies the factual accuracy of content generated by large language models (LLMs).
-The fact-checking is conducted using the HELM dataset on the final trained models (1B SeND and 1B Normal Training).
-A higher FactScore indicates improved factual precision.
-When evaluated on 100 data points from the HELM dataset, the 1B SeND model achieves a FactScore of 0.07, whereas the 1B Normal Training model attains 0.05, demonstrating a 40% improvement in factual accuracy, even during test time.
-This highlights the efficacy of SeND in enhancing the factual certainty of the model.
-Note that SeND is not a replacement for post-hoc methods like RAG (Gao et al., 2024b), but rather to complement them.
-The robustness of this training protocol is essential as we aim for it to be applied across many different fields.
-In light of this, we present the results of finetuning on MedHALT 2k with and without SeND in Figure 4(b).
-We observe a similar trend in 4(b) as shown in Figure 4(a), where standard finetuning increases the EES score throughout training, showing that the model is not taking into account hallucinations and factuality during its training.
-In Figure 4(b), there is an improvement in the trajectory of EES as training continues, showing that our model is in fact able to incorporate factuality as a metric to account for during training of the model.
-Although the decrease in EES is marginal for MedHALT 2k, we can attribute this to the dataset it is being trained on.
-The relatively small difference observed between the training protocol behaviours could be due to MedHALT 2k data never being seen before the finetuning phase whereas HELM data has been seen.
-In this case, it may be beneficial to delay the onset of SeND, as high variability between checkpoints on new training instances is expected.
+To evaluate SenD, we use Pythia 1B model (Biderman et al., 2023), Llama 3.2 1B, and Llama 3.1 8B (Dubey et al., 2024) continuing their training on specific datasets rather than restarting pretraining for efficiency.
+We continually train the models on the following datasets: HELM, consisting of Wikipedia text (Su et al., 2024), MedHALT, a medical dataset emulating real-world entrance exam questions (Pal et al., 2023), and LegalBench consisting of data for reasoning in LLMs (Guha et al., 2023).
+Note that HELM and MedHALT are specifically designed for hallucination detection/mitigation in LLMs.
+SenD is trained on all datasets using 200 data points and and 2,000 points (referred to as 2k) for MedHALT due to the medical domain importance.
+SenD implements the EigenScore reduction technique from Section 3.2 and detects SEIs using a 3- checkpoint window on a specialized hallucination tracking dataset.
+The distance between checkpoints and the dropout rate are tunable hyperparameters.
+Given our ablation study in Appendix C, we opt for and Threshold=3 for the experiments.
+SEIs in the penultimate layer are identified based on their variability across checkpoints and are deterministically dropped for the subsequent 3 training checkpoints.
+This is repeated at each 3- checkpoint interval until loss convergence, effectively mitigating hallucination tendencies and oscillations.
+4.2 Performance of SenD on Pythia and Llama Models
+Pythia and Llama training results are illustrated in Figure 4.
+To validate that EES accurately approximates the EigenScore metric; we compare the model’s progress during training detailed in Appendix B.6.
+Upon confirming that, we proceed to compare the performance of Pythia 1B, Llama 3.2 1B, and Llama 3.1 8B trained using training without dropout to that of SenD (Figure 4).
+As shown in the figure and detailed in Appendix D, across all three models and most domains, training with SenD results in a greater reduction in EES during training.
+In most cases, the final model trained with SenD achieves a lower EES compared to standard training which increases the EES score, demonstrating its effectiveness in reducing hallucination variance and improving the model’s final certainty.
+In addition, there seems to be better improvements using SenD in LegalBench settings.
+A possible explanation for SenD’s superior performance on LegalBench compared to the other two domains which are more general is that early in training, the model sees less domain-specific data, leading to better generalization.
+However, confirming this hypothesis requires further investigation beyond the scope of this work.
+To assess the effectiveness of SenD in comparison to other state-of-the-art factuality metrics, and not solely rely on EES for hallucination detection, , we employ the FactScore metric (Min et al., 2023), which quantifies the factual accuracy of content generated by large language models (LLMs) and HaluEval in Summarization setting (Li et al., 2023).
+The fact-checking for FactScore is conducted using the HELM dataset on the final trained models (Pythia 1B SenD and 1B Normal Training).
+A higher FactScore indicates improved factual precision and for HaluEval, higher Exact Match, Accuracy, and Correctness imply better matching with ground truth facts.
+FactScore is done on 100 and 1000 data points subsequently and table 1 presents the results which imply the better performance of the end model trained with SenD up to 40%.
+This also depicts the correlation between low EES and high FactScore or HaluEval metrics which are other SOTA metrics for hallucination detection.
+| Task | Metric | SenD | Normal |
+| HaluEval Summarization (LMEval) | Accuracy | 0.016 | 0.014 |
+| Correctness | 0.027 | 0.027 | |
+| Exact Match | 0.589 | 0.496 | |
+| FactScore (100 points) | Score | 0.07 | 0.05 |
+| FactScore (1000 points) | Score | 0.08 | 0.06 |
+Since SenD is the first method to focus on Hallucinations during the training of LLMs, there are no baselines or SOTA methods to compare it to.
+However, one could treat SenD as a post-hoc method and compare it to RAG (Lewis et al., 2021) to get a sense of how different training based methods compare to in-context methods.
+When tested on 1000 data points, SenD-continually-trained Pythia achieves a FactScore of 0.07 as in Table 1, while the base Pythia with RAG scores 0.25, highlighting RAG’s effectiveness in reducing hallucinations by providing context.
+However, to make the comparison fair, applying RAG to a SenD-trained model achieves a higher FactScore of 0.28, outperforming RAG on a normally trained model by 12%.
+This indicates that even though SenD is not meant to outperform post-hoc methods, SeND combined with RAG enhances the end model’s factual performance compared to RAG on a normally trained model.
 5 Conclusion & Future Work
@@
 (2024).
-We showed empirically that our Sensitive Neuron detection method significantly reduces the EigenScore on inference of LLMs throughout various stages of training (Figure 2).
-Following the success of the Sensitive Neuron method, we moved on to the application of a hallucination reduction method on training.
-We show through finetuning that we are able to fix the oscillatory behaviour initially seen throughout training and reduce the EES of finetuned models as shown in Figure 4 by modifying the internal mechanics of training with SEnsitive Neuron Dropout.
-At test time we achieve a 40% increase in FactScore performance, verifying that SeND provides a substantial improvement to current training protocols.
+We showed empirically that our SEI detection method significantly reduces the EigenScore on inference of LLMs throughout various stages of training (Figure 2).
+Following the success of the SEI method, we moved on to the application of a hallucination reduction method on training on Pythia and Llama models and various domains.
+We show through training with SenD that we are able to fix the oscillatory behaviour initially seen throughout training and reduce the EES of finetuned models as shown in Figure 4 by modifying the internal mechanics of training with Sensitivity Dropout.
+At test time we achieve a 40% increase in FactScore performance and improvement of other SOTA hallucination detection metrics, verifying that SenD provides a substantial improvement to current training protocols both during and after training.
+Note that although SenD has only been applied to finetuning training in this paper, the training framework is applicable to all stages of training.
+We encourage future work to implement SenD on larger training sets, such as pretraining, to see how SenD performs in these environments.
 To further advance our work, we plan to scale SenD to larger datasets and models, as current experiments were limited by compute constraints with larger LLMs.
-Demonstrating SeND’s effectiveness on larger open-source models like Meta’s LLaMA 3.1 (Dubey et al., 2024) will provide crucial evidence for organizations developing state-of-the-art LLMs to incorporate SeND into their training protocols, ultimately improving model safety.
+Demonstrating SenD’s effectiveness on larger open-source models like Meta’s LLaMA 3.2 405B (Dubey et al., 2024) will provide crucial evidence for organizations developing state-of-the-art LLMs to incorporate SenD into their training protocols, ultimately improving model safety.
 Given that SenD targets variance reduction during training, we anticipate even greater gains on larger LLMs, where the higher inherent variance may amplify the regularization effect and lead to more significant improvements.
 Appendix A Additional Experiments
-A.1 Drastic Embedding Changes leading to Sensitive Neurons
+A.1 Drastic Embedding Changes leading to Sensitive Embedding Indices
 Looking at internal states of the model allows getting a deeper understanding of the dynamics that could be leading to the oscillatory behaviour seen in Figure 1.
@@
 It can be observed that there were specific embedding activations that experienced drastically more change relative to the rest of the embeddings.
-This is the main source of motivation to further define Sensitive Neurons (Definition 3.3).
+This is the main source of motivation to further define SEIs (Definition 3.3).
+A.2 Hallucination Oscillations Across Model Sizes
+Figures 6, 7, and 8 show our study of hallucination oscillations during the training of Pythia models.
+An overall observation across the plots is that as opposed to our intuitive expectation which is a linear decrease of the hallucination detection metric when the model scales linearly, neither the oscillations during the training of the model decrease, nor the end model reaches its optimal state in terms of the hallucination metric.
 Appendix B Efficient EigenScore (EES) Derivation
@@
 B.6 EigenScore and EES training trajectories
-To demonstrate that our EigenScore approximation method, EES, is a good metric, we record the progress of Pythia 1B finetuning on the HELM dataset using both EigenScore and EES hallucination performance metrics (Figure 7).
+To demonstrate the high correlation between EigenScore and EES, we record the progress of Pythia 1B finetuning on the HELM dataset using both EigenScore and EES hallucination metrics (Figure 10).
 Albeit a different scale and window, the trajectories, magnitude and shape of the graphs are nearly identical while EES takes only 4 minutes to calculate and EigenScore takes approximately 8, an astounding 2x increase in compute speed.
 These results show that our metric closely resembles the target metric while greatly reducing the required computational resources.
+Appendix C Ablation study on K and Step Thresholding for SenD
+Figure 11 shows the ablation study done on and Figure 12 illustrates the ablations study done on the Step Threshold for SenD experiments.
+As depicted, and Threshold = 3 are chosen for our experiments except for Llama 3.1 8B model which due to its larger size requires more embedding indices to be dropped, hence adapting to .
+Appendix D Additional Pythia 1B, Llama 3.2 1B, and Llama 3.1 8B Training with and without SenD
+Here, we present additional experimental results of training Pythia and LLaMA on multiple domains.
+Figure 14(b) supplements the results discussed in Section 4 by illustrating the training procedures on the Medical domain.
+In the Pythia 1B setting, the EES achieved with training using SeND remains consistently lower than that of normal training and exhibits fewer oscillations throughout the training process.
+In the LLaMA 3.1 8B setting, while both approaches show an increase in the EES metric during training, the final model trained with SeND achieves a lower EES, indicating a reduced likelihood of hallucinations in this domain.
+Figure 14 visualizes the training results for the LLaMA 1B model across all domains.
+Similar trends to the other experimental settings are observed.
+In the HELM and LegalBench settings, the EES of the model trained with SeND shows a significant reduction by the end of the training.
+However, in the MedHalt setting, this reduction is less pronounced by the conclusion of training.
+Figure 15 depicts the standard deviations of 5 runs experimented on the Pythia 1B model.
+As shown, the error shading area is not too wide hence proving the reliability of the results.